THE DISSOCIATIVE RECOMBINATION OF CH_{3}O^{+} AND CD_{3}O^{+} CATIONS

\underline{MATHIAS HAMBERG}, Roslagstullsbacken 5:310 ZIP 114 21 Stockholm Sweden.

Reactions with the methoxy radical CH_{3}O^{+} have found to be significant pathways of production and destruction of important interstellar molecules and ions like NH_{4}^{+}, NH_{3}, H_{2}O, H_{3}O^{+} ^{a}. These processes compete with the dissociative recombination of the named ions. Therefore, measurements of the rates, cross-sections and branching ratios of the dissociative recombination of the iosotopomers CH_{3}O^{+} and CD_{3}O^{+} have been perfomed at the CRYRING storage ring located at Stockholm, Sweden. Preliminary evaluation of the data yielded a reaction rate coefficient of 6.8 ∗ 10^{−7} (\frac{T}{300})^{−0.63} cm^{3}mol^{−1}s^{−1} and 8.4 ∗

10^{−7} (\frac{T}{300})^{−0.62} cm^{3}mol^{−1}s^{−1} for CH_{3}O^{+} and CD_{3}O^{+}, respectively. Calculation of the branching ratios rendered the following branching ratios for the reaction channels leading to the fissure of the C-O bond:, H_{2}O+CH (0.7%), CH_{2} +OH (1.6%); D_{2}O + CD (2%), CD_{2} + OD (6%) respectively. The rest of the dissociative recombination rates was contributed from channels leaving the C-O bond intact: 92% and 98% for CH_{3}O^{+} and CD_{3}O^{+}, respectively. Contrary to the CH_{3}OH_{2}^{+} ion^{b}, the pathways preserving the bond between the heavy atoms dominate.

aS.A. Haider, Anil Bhardwaj, ICARUS, 177, 196.

bGeppert, W. D. Semaniak, J., Hellberg, F, Österdahl, F., Roberts, H.,

Millar, T. J, Hamberg, M., Thomas, R. D., af Ugglas, M., Ehlerding, A.,

Zhaunerchyk, V., Kaminska, M., Larsson, M., Faraday Discussions, 133, 2006,

accepted for publication.